OpenStudy (anonymous):

If f and g are differentiable functions for all real values of x such that f(1) = 4, g(1) = 3, f '(3) = −5, f '(1) = −4, g '(1) = −3, g '(3) = 2, then find h '(1) if h(x) = the quotient of f of x and g of x..

OpenStudy (anonymous):

\[h(x)=\frac{f(x)}{g(x)}\]?

OpenStudy (anonymous):

\[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] you are given each number, plug them in

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The answer is not listed